Factorize: x^6 - 10x^3 + 27:(^-raised to)

Grade:10

1 Answers

Askiitians Expert Mohit Singla
14 years ago

 

Dear Ajinkya

x6 - 10x3 + 27=0

 

let x3 =t

so t2 -10t +27 =0

discriminant =(-10)^2-4(1)(27)  =-8 which is less than 0.

So,discriminant of this equation is negative so  no real value of t.

So complex roots exist for this equation,which are {10+i(8)^(1/2)}/2 and {10-i(8)^(1/2)}/2.

So this equation has no real value of x.


Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
MOHIT